OpenStudy (anonymous):

Find the volume of the solid generated by revolving the region bounded by the graphs of y=3sqrt(x) and y=(x^2/8) around the x-axis in cubic units rounded to the nearest hundredth.

OpenStudy (anonymous):

The answer is 587.2un^3, but no matter what I try, I cannot get to that value. Help with figuring out the steps would be much appreciated.

OpenStudy (lgbasallote):

a usual mistake of people is limits of the integral...or the method used (i.e. shell, washer, disk, etc)

OpenStudy (zzr0ck3r):

are you doing double integral or washer disk...?

OpenStudy (anonymous):

I have tried disk, washer, and shell methods but can't come up with that number. I think it should be the washer method, but I'm doing something wrong in the math.

OpenStudy (kainui):

First off, what does that shape look like?  Now what are your limits of integration?

OpenStudy (lgbasallote):

it seems the washer is the most preferrable according to @Kainui 's graph

OpenStudy (anonymous):

do you mean the values that define the integral? I was using 0 and the point where they graphs intersect at 8,6535

OpenStudy (anonymous):

But now that I've seen a bigger picture than my calculator, it looks like 0 is wrong.

OpenStudy (anonymous):

^^that was 8.6535

OpenStudy (kainui):

Wait, is that \[\frac{ x^2 }{ 8 } =or =x^\frac{ 2 }{ 8 }\]

OpenStudy (kainui):

Oh I see ok lol

OpenStudy (anonymous):

The first equation.

OpenStudy (kainui):

Alright, so you know the limits of integration, good. Now what shapes are you making with those two graphs and how?

OpenStudy (kainui):

I like to think of it this way. You're making two solid shapes but you're subtracting one of the solid shapes from the other.

OpenStudy (anonymous):

Okay, so we're still using 0 and 8.6535 for the region correct? And the way I would set it up is V=pi\[\int\limits_{0}^{8.6535}\]((3\[\sqrt{x)}\]-(x^2)/8

OpenStudy (anonymous):

Sorry I couldn't really figure out the equation machine. so integrand would be (3sqrt(x))^2-((x^2)/8)^2

OpenStudy (kainui):

Well I just checked and I found that they intersect at about 8.320335

OpenStudy (anonymous):

Darn it I was reading the wrong value. 8.6535 is the y value, not the x. Let me try plugging in 8.32035

OpenStudy (anonymous):

With the region (0,8.32035) I got the answer 186.9155

OpenStudy (kainui):

\[\pi[ \int\limits_{0}^{8.32}3\sqrt{x}dx-\int\limits_{0}^{8.32}\frac{ x^2dx }{ 8 }]\] that's the integration I found, I might have made a mistake I didn't really think about it too much

OpenStudy (kainui):

Yeah I forgot to square the function lol

OpenStudy (kainui):

Does that look like what you were doing? I couldn't really tell, but I can explain what I did and why.

OpenStudy (anonymous):

Yes, but I squared both of the functions. So after squaring and integrating I got pi[((9/2)x^2)-((x^5)/320)]

OpenStudy (kainui):

Yeah that's what I got, I'm about to plug in the answer myself.

OpenStudy (kainui):

I'm also getting 186.915. How do you know that's not the right answer lol?

OpenStudy (anonymous):

Haha it is a practice version of the final exam which came with an answer key - so hopefully the key would be correct...

OpenStudy (kainui):

Oh wait.

OpenStudy (kainui):

What's 186.92 times pi?

OpenStudy (kainui):

;)

OpenStudy (anonymous):

Oh jeeze. Lol now I feel really silly. Thanks for walking the whole thing through with me though!

OpenStudy (kainui):

Haha no problem, I made the same mistake too rofl.